# Shower formation in the presence of a string-inspired foam in space-time

### v1

@@ 1 Introduction
We know that Lorentz symmetry is crucial in fundamental physics and any idea and/or experimental indication of its violation which might characterize some quantum gravity (QG) theories [1], should be treated with care. Models entailing Lorentz violation (LV) generically predict energy-dependent speed variations for relativistic probes provided the space-time probed at a small enough scale could appear complicated—something akin in complexity to a turbulent froth that Wheeler dubbed space-time foam [2]. Energetic $\gamma$-rays from cosmic sources like gamma-ray bursts (GRBs) provide a window to search for such effects through the so-induced time delays [3, 4]. Specifically, the authors of [5, 6] analyze a dataset consisting of 14 multi-GeV photons from 8 Fermi GRBs and upon incorporating a constant intrinsic lag at the source (see also [7, 8, 9, 10, 11, 12, 13, 14]) in the observed time lags, they suggest a light-speed variation, $v(E)=c(1-E/E_{\textrm{LV}})$, with the LV scale, $E_{\textrm{LV}}\sim 3\times 10^{17}$ GeV. This corresponds to a subluminal dispersion relation,
@@
with the sign factor $s=+1$ and $E_{\textrm{LV}}$ approaching the order of Planck-scale $E_{\textrm{Pl}}\sim 10^{19}$ GeV. We shall use the units so that $c=\hslash=1$ with $c$ the constant speed of light. Recently, the Monte Carlo (MC) analysis [15] implies the necessity of introducing energy-dependent intrinsic delays [16] in order for 3 more remarkable events across GeV to TeV bands to be robustly fitted within the same physics scenario (1[17]. With the full Bayesian parameter estimation strategy employed, a recent work in [18] finds that the null hypothesis of dispersion-free vacuum $E=k$, or, equivalently, the constant light speed $v=1$ is rejected at a significance level of $3.1\sigma$ or higher.
While one might consider earlier studies [7, 8, 9, 5, 6, 10, 11, 12, 13, 14] not so reliable (due to imperfect knowledge of the radiation mechanism of GRBs) and LV hypothesis too premature the statistical analysis involving MC simulation has indicated the robustness of the finding. Besides conventional astrophysical explanations like matter effects which can mimic LV [8], some exotic interpretations of the result were recently suggested [19, 20, 21]. These are based on string models [22, 23, 24, 25, 26, 27, 21, 28, 29, 30] of space-time foam-induced refractive index in vacuo $\eta$ which is subluminal, growing linearly with photon energy $\omega(=E)$ and, suppressed by a single power of the string mass scale, $M_{s}$, viz.,
@@
Here the functional form of QG scale $f(M_{s})\eqqcolon M$ depends on foam types and $M\sim E_{\textrm{LV}}$, indicated by the above phenomenological finding. The important point to note here is that such explanations are also consistent with other stringent tests of LV, in particular constraints from $\gamma$-decay and birefringence effect (for details see [19]).
@@
The common wisdom regarding the impact of LV is that, although it affects light propagation in free space (and possibly the thresholds of photoproduction collisions with soft photon of cosmic background lights) the effect of the modified dispersion relation on interactions relevant for the detection of high-energy emissions is negligible. Nonetheless, in [31], the author first points this may not hold true: electromagnetic (EM) interactions like pair creation in the detectors onboard satellites or Cherenkov process in ground-based observatories may be modified by LV. A later analysis further notes [32] that dispersion effect, of the type (1), will imply unacceptable changes in the development of EM shower cascade. In particular, a strong suppression of photon “decay” into an $e^{\mp}$ pair in the Coulomb field of nuclei, the process mainly responsible for the formation of photon-initiated atmospheric showers [33], could be expected [34]. For PeV $\gamma$-rays at the Large High Altitude Air Shower Observatory (LHAASO) [35], showers that are observed with no deviation from standard cascade theory have been used recently in [36] to severely constrain the field theory of [37]: $E_{\textrm{LV}}\gtrsim{\cal O}(10^{20})$ GeV for subluminal LV photons (1). However, all of these conclusions are based on assuming energy conservation which may be invalid in the above-mentioned string-inspired “foamy” situations.
@@
Thus, in this Letter we aim to discuss the influence from stringy space-time foam and the associated energy fluctuations on particle shower formation, which is still unknown, and further examine whether the obtained results from the phenomenological or, at most, the effective field theory approaches can still apply, once such fluctuations induced by the foam are present. Here, it is worth noticing that different formulations in this framework have suggested different modifications of the standard physics. For example, for the foam type in [26], which does not entail deformed relations between energy and momentum, there is no mechanism for energy nonconservation despite the presence of subluminal photons. In this Letter, however for our purposes, we only focus on discussing QG effect originating from an adaption of the model recently formulated in our Ref. [21] on shower development. It is shown that under certain conditions the formation length for pair creation and bremsstrahlung and thus qualitatively, their cross-section, are unaffected; then the light-speed variation would not lead to drastic changes to photon showers in the context of such models and evades the above constraints.
@@ 2 String foam and the shower formation
In string models for space-time foam, the in-vacuo propagation of photons is modified as initially inspired from Liouville QG [38], with their travel times delayed by amounts proportional to $\eta-1$ (c.f. (2)), due to D-brane space-time defects. These defects break Poincaré invariance, and their recoil during interactions with an open string representing a photon breaks local Lorentz invariance. They give space-time a foamy nature so the situations were termed D-foam. The cosmological setup of D-foam is based on braneworlds moving in a higher-dimensional bulk space that contains a “gas” of D0-branes (D-particles) [24] of type I/IIA strings. We consider it to be present in models of phenomenological interest [22, 23, 24, 25, 26, 27, 19, 20, 21, 28, 29] as even if the theory admits no D-particles one has effective ones [30]. As our brane roams in the bulk, they cross it randomly, and collide with the photon resided on it and it is not possible in general to capture the process with local effective Lagrangians.
We have previously investigated stochastic effects of the foam in a model of this kind in [21] which will violate CPT, and neutrinos may experience a so-induced CPT violation during propagation in this QG “medium”, dictated by the dispersion relations:
@@
In order to discuss cosmic photon-induced showers let us first derive the photon dispersions, via considering the nontrivial momentum transfer $\Delta k_{i}\equiv rk_{i}$, in each collision of a photon off a D-(particle)-defect, which recoils at a velocity, $u$ with its components along the brane, $u_{i}=rg_{s}k_{i}/M_{s}$ [22, 23, 24]. We assume isotropic random recoil, as in [21], i.e., the momentum fraction variable $r$ vanishes on average $\langle\!\langle r\rangle\!\rangle=0$ and $\langle\!\langle r^{2}\rangle\!\rangle\eqqcolon\sigma^{2}<1$, where $\langle\!\langle\cdot\cdot\cdot\rangle\!\rangle$’s are statistical averages over a collection of quantum-fluctuating defects met by the photon. Here, the variance $\sigma$ plays a role similar in nature to the parameter $\zeta$ for neutrino species ($\sigma\neq\zeta$ in general). [^1] The topologically nontrivial effects of the foam on particle propagation do not obey the principle of equivalence [39, 40], in the sense of being universal for all types of excitations of the Standard Model; moreover for reasons of charge conservation such a medium of defects is “transparent” for charged particles, such as electrons (as indicated by experiment [41]). The metric surrounding the struck defect is of Finsler type $G_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$ and $h_{0i}=u_{i}$ modifying the energy of the photon via $k^{\mu}k^{\nu}G_{\mu\nu}=0$. One obtains after averaging over foam particles the modified energy, $\omega\simeq k[1+\frac{1}{2}\sigma^{2}k^{2}/M_{D}^{2}]$, used in the kinematics: $\omega=\omega^{\prime}+\langle\!\langle M_{D}u^{2}/2\rangle\!\rangle$, where $\frac{M_{s}}{2g_{s}}u^{2}$ is the kinetic energy of heavy defects in the case of nonrelativistic recoil, $u_{i}\ll 1$. The final dispersion relation found from $\omega^{\prime}$ in the leading approximation yields,
@@
here $\sigma^{2}$ a free parameter, depends on foam density $n_{D}$ (and microscopic details that cannot be determined quantitively presently). It is clear that this subluminal dispersion effect leads, through the appropriate group velocity ($v\coloneqq\partial\omega/\partial k$ assumed to hold), to a QG refractive index (2) for photons: $v\simeq 1-\omega\sigma^{2}/M_{D}$, of the kind favored [19] by current time-lag studies of GRB photons [7, 8, 9, 5, 6, 10, 11, 12, 13, 14, 16, 18, 17, 15]. For a uniform D-defect density, at least for late eras of the Universe history where $\sigma$ is constant, an order of magnitude $\sigma^{2}\sim{\cal O}(1)$ is sufficient to reproduce the finding of light-speed variation mentioned earlier, assuming a Planckian D-particle mass $\sim 10^{18}$ GeV. In cases of lower $M_{D}$ the required $\sigma^{2}\ll 1$ is easily satisfied with more dilute foam populations.
@@
It has been argued [36] however that, due to the presence of the modified photon dispersion (1) which, as we see, may originate from a string-inspired scenario (c.f. (4)) the cross-sections of standard (perturbative) processes are modified. In particular, any allowed interaction channel is suppressed for subluminal photons, since such LV photons lack energy compared to the momentum. Earlier as argued in [32] they may even suppress the cascade generated by cosmic $\gamma$-rays, causing the showers to be formed deeper in the atmosphere and in detectors. The absence of the effect in the measured spectrum sets restrictive (shower formation) constraints on $E_{\textrm{LV}}$ [36]. However, these considerations also concern strict energy–momentum conservation (assumed in prior works), not only modified dispersion relation.
@@
In general, in the space-time foam picture energy is conserved only in a statistical sense (see the review [42]), however, even this can be violated due to unobserved degrees of freedom (from the point of view of a low-energy braneworld observer) associated with the D-particle recoil in the model of [21]. Indeed, there are quantum fluctuations $\delta E_{D}$ in the total energy in particle reactions [23]. For a three-body decay process in the recoil foam adapted to the situation here discussed, the energy of the incoming particle $E$ fluctuates, due to the presence of stochastic D-foam, by the amount of $(\varsigma_{I}/M_{D})p^{2}$ where $p$, the momentum of the particle emitted during the decay of a particle of effective energy $E+\delta E_{D}$. The proportional factor $\varsigma_{I}$ expressing the intensity of such fluctuations can vary independently of the foam dispersion factor $\sigma$ and values $\varsigma_{I}>0$, as in principle permitted by the theory, will be seriously considered here. Now we are going to see how this affects the conclusion drawn from other LV approaches [32, 37], yielding unacceptable anomalies in the formation of cosmic $\gamma$-ray showers. As we shall see, the peculiarity of this foam in resulting in such energy violations is crucial for accommodating the result of light-speed variation, without conflicting with the constraints from shower measurements.
@@
At the Earth photons initiate almost purely EM showers via pair creation and bremsstrahlung [33]. Additional processes are important only at highest energy ($10^{19}-10^{20}$ eV) so for GeV to sub-PeV $\gamma$-rays we ignore them for simplicity except where noted. Lacking at present a complete theory of the matter/D-foam interactions we do not deal with the cross-sections directly and shall restrict ourselves to calculate the length $\lambda_{f}$, over which the processes are considered taking place according to uncertainty principle, that is formation (or coherent) length [43].
@@
where $p$, $p^{\prime}$ indicate the outgoing momenta of electron and positron and photon’s motion is in $z$ axis, with its momentum $k$. The nucleus’ Coulomb field has a larger component perpendicular to $z$ axis, leading to a transverse momentum kick $\Delta q_{\perp}\sim{\cal O}(m_{e})$ ($m_{e}$, the electron rest mass), while the longitudinal momentum $\Delta q_{z}$ is small, $\Delta q_{z}\ll k,\Delta q_{\perp}$ [34], due to the photon’s high energy $\omega\gg 2m_{e}$. For this regime one can ignore collision angles. In the presence of string D-foam, one should make use of the dispersion including the foam induced LV (4). For electron (positron) its law of dispersion is preserved as noted, for gauge invariance [39, 40], so we have $E^{2}-m_{e}^{2}=p^{2}$ and then,
@@
We come to discuss the formation length, $\lambda_{f}$, dictated by Heisenberg’s uncertainty principle. While the latter suffers a modification, due to the existence of the minimal (string) length $\ell_{s}=1/M_{s}$ in target space-time [44],
@@
where $\langle X_{0}\rangle\simeq 50~\textrm{g}\cdot\textrm{cm}^{-2}$ with $\sigma_{\textrm{BH}}\approx 0.51$ b in air. Then, the probability for a $\gamma$-ray to develop a shower and then be detected as penetrating through the Earth’s atmosphere is given in [36] and when adapted to our case reads
@@
It is from these considerations that interesting constraints on subluminal-photon linear LV were derived recently from the LHAASO data for TeV–PeV photon-induced air showers [35]. Unlike the linear LV model of [32, 37], however, in our case, such limits are not imposed on dispersion parameter $\sigma^{2}$ which is related to $E_{\textrm{LV}}$ in Eq. (1) but for $2\sigma^{2}-\varsigma_{I}$. Naively, one could use the result they derive to make an order of magnitude estimate in the parameters of the present model: [^2] We notice that for the result of [36], the effect of LV-modified pair creation on soft cosmic background photons leading to the increase of the flux is also taken into account, whilst it is subleading compared to the shower formation effects (also for priorly studied spectra [45, 46]). However, the existence of such modifications in the present model has not been clarified, hence temporally ignored here and is left for future investigation.
@@
The other effect that opposes those from general and/or field-theoretic description of subluminal LV photons is that if $\varsigma_{I}>2\sigma^{2}$, Eq. (2) turns into enhancement which implies that, the formation length increases instead of being shortened. And the factor $S$ even becomes infinite at the critical energy $\omega_{\textrm{cr}}=\bigl{[}\frac{8M_{D}m_{e}^{2}}{(\varsigma_{I}-2\sigma^{2})}\bigr{]}^{1/3}$ and negative above it, while the situation could be compensated by suppression factors associated with interactions that can disrupt coherence over such a longer length. The effects, that increase $\Delta q_{z}$ and reduce $\lambda_{f}$, include, for example, the Landau–Pomeranchuk–Migdal effect (see, e.g., [43, 33]) due to multiple scattering centers. The so-reduced BH cross-section delays the development of the EM cascade initiated by a photon competing with foam effects which decrease now $\langle X_{0}\rangle$ and $\langle X_{\max}\rangle$. It is worth noting that the length of shower development $\Delta X$ shifting $X_{\max}$ as $X_{\max}=X_{0}+\Delta X$ is assumed unmodified in the above discussion, since the secondary interactions in the cascade are less energetic than the primary one. While, in principle, it may be affected by “exotic” physics studied here, not accounting for this leaves the estimate (19) more conservative.
@@
where $p$ and $p^{\prime}$ stand now for the electron momenta before and after the interaction, and the photon emission angle is neglected for highly relativistic electron $E\gg m_{e}$. One difference herein is that, unlike pair creation where the energy partition in the produced $ee^{+}$ pair is most likely symmetric as we assumed above, [^3] It was shown [34] that LV may enhance the leading particle effect as the cross-section is peaked at $E\sim\omega$ and $\omega-E\sim\omega$. But, there is no good reason to expect the same for D-foam cases where, unlike the model studied therein, to first order in $1/M_{s}$, the matrix element for any process agrees with the one in the special-relativistic field theory so we consider symmetric pairs as an illustration. the probability of asymmetric share is significant in bremsstrahlung. The resulting spectrum is dominated by photons emitted at very low energy ($\omega\ll E$) and diverges as $\omega\rightarrow 0$ ($1/\omega$ dependence). Then the effects of foam for such configurations of soft bremsstrahlung photons are negligible, compared to the first two conventional, leading terms as in (8):
@@
While for photons at moderate energies, the process may be mildly affected, nevertheless it is important to note that in order for an electron to emit a real photon as it interacts with the field of the nucleus, part of its surrounding virtual photon field shakes loose, a case for which no modification is induced in this scenario. As we have explained, LV effect within such D-brane models is due to multiple interactions between photons and D-particles, hence cannot be relevant in any mere process of photon production. This is another difference from the phenomenological consideration of [32], where due to electron LV effect, bremsstrahlung is strongly suppressed, similarly to pair creation on nuclei by photons, reducing the formation length $\lambda_{f}$ from (21), and hence the probability for radiation. There is currently no direct limit from bremsstrahlung though as the latter is usually viewed as unmodified, for practical (LV constraining) purposes, as also mentioned before.

### v2

@@ 1 Introduction
We know that Lorentz symmetry is crucial in fundamental physics and any idea and/or experimental indication of its violation which might characterize some quantum gravity (QG) theories Li:2025stf , should be treated with care. Models entailing Lorentz violation (LV) generically predict energy-dependent speed variations for relativistic probes provided the space-time probed at a small enough scale could appear complicated—something akin in complexity to a turbulent froth that Wheeler dubbed space-time foam Wheeler:1998stf . Energetic $\gamma$-rays from cosmic sources like gamma-ray bursts (GRBs) provide a window to search for such effects through the so-induced time delays Amelino-Camelia:1997ieq , Jacob:2008bw . Specifically, the authors of Xu:2016zxi , Xu:2016zsa analyze a dataset consisting of 14 multi-GeV photons from 8 Fermi GRBs and upon incorporating a constant intrinsic lag at the source (see also Shao:2009bv , Zhang:2014wpb , Amelino-Camelia:2016ohi , Xu:2018ien , Liu:2018qrg , Chen:2019avc , Zhu:2021pml , Zhu:2022usw ) in the observed time lags, they suggest a light-speed variation, $v(E)=c(1-E/E_{\textrm{LV}})$, with the LV scale, $E_{\textrm{LV}}\sim 3\times 10^{17}$ GeV. This corresponds to a subluminal dispersion relation,
@@
with the sign factor $s=+1$ and $E_{\textrm{LV}}$ approaching the order of Planck-scale $E_{\textrm{Pl}}\sim 10^{19}$ GeV. We shall use the units so that $c=\hslash=1$ with $c$ the constant speed of light. Recently, the Monte Carlo (MC) analysis Song:2025akr implies the necessity of introducing energy-dependent intrinsic delays Song:2024and in order for 3 more remarkable events across GeV to TeV bands to be robustly fitted within the same physics scenario (1Song:2025qej . With the full Bayesian parameter estimation strategy employed, a recent work in Song:2025myx finds that the null hypothesis of dispersion-free vacuum $E=k$, or, equivalently, the constant light speed $v=1$ is rejected at a significance level of $3.1\sigma$ or higher.
While one might consider earlier studies Shao:2009bv , Zhang:2014wpb , Amelino-Camelia:2016ohi , Xu:2016zxi , Xu:2016zsa , Xu:2018ien , Liu:2018qrg , Chen:2019avc , Zhu:2021pml , Zhu:2022usw not so reliable (due to imperfect knowledge of the radiation mechanism of GRBs) and LV hypothesis too premature the statistical analysis involving MC simulation has indicated the robustness of the finding. Besides conventional astrophysical explanations like matter effects which can mimic LV Zhang:2014wpb , some exotic interpretations of the result were recently suggested Li:2021gah , Li:2021tcw , Li:2022sgs . These are based on string models Ellis:1999uh , Ellis:2000sf , Ellis:2004ay , Mavromatos:2005bu , Ellis:2008gg , Mavromatos:2012ii , Li:2022sgs , Li:2023wlo , Li:2024crc , Li:2009tt of space-time foam-induced refractive index in vacuo $\eta$ which is subluminal, growing linearly with photon energy $\omega(=E)$ and, suppressed by a single power of the string mass scale, $M_{s}$, viz.,
@@
Here the functional form of QG scale $f(M_{s})\eqqcolon M$ depends on foam types and $M\sim E_{\textrm{LV}}$, indicated by the above phenomenological finding. The important point to note here is that such explanations are also consistent with other stringent tests of LV, in particular constraints from $\gamma$-decay and birefringence effect (for details see Li:2021gah ).
@@
The common wisdom regarding the impact of LV is that, although it affects light propagation in free space (and possibly the thresholds of photoproduction collisions with soft photon of cosmic background lights) the effect of the modified dispersion relation on interactions relevant for the detection of high-energy emissions is negligible. Nonetheless, in Kifune:1999ex , the author first points this may not hold true: electromagnetic (EM) interactions like pair creation in the detectors onboard satellites or Cherenkov process in ground-based observatories may be modified by LV. A later analysis further notes Vankov:2002gt that dispersion effect, of the type (1), will imply unacceptable changes in the development of EM shower cascade. In particular, a strong suppression of photon “decay” into an $e^{\mp}$ pair in the Coulomb field of nuclei, the process mainly responsible for the formation of photon-initiated atmospheric showers Risse:2007sd , could be expected Rubtsov:2012kb . For PeV $\gamma$-rays at the Large High Altitude Air Shower Observatory (LHAASO) LHAASO:2021gok , showers that are observed with no deviation from standard cascade theory have been used recently in Satunin:2023yvj to severely constrain the field theory of Myers:2003fd : $E_{\textrm{LV}}\gtrsim{\cal O}(10^{20})$ GeV for subluminal LV photons (1). However, all of these conclusions are based on assuming energy conservation which may be invalid in the above-mentioned string-inspired “foamy” situations.
@@
Thus, in this Letter we aim to discuss the influence from stringy space-time foam and the associated energy fluctuations on particle shower formation, which is still unknown, and further examine whether the obtained results from the phenomenological or, at most, the effective field theory approaches can still apply, once such fluctuations induced by the foam are present. Here, it is worth noticing that different formulations in this framework have suggested different modifications of the standard physics. For example, for the foam type in Ellis:2008gg , which does not entail deformed relations between energy and momentum, there is no mechanism for energy nonconservation despite the presence of subluminal photons. In this Letter, however for our purposes, we only focus on discussing QG effect originating from an adaption of the model recently formulated in our Ref. Li:2022sgs on shower development. It is shown that under certain conditions the formation length for pair creation and bremsstrahlung and thus qualitatively, their cross-section, are unaffected; then the light-speed variation would not lead to drastic changes to photon showers in the context of such models and evades the above constraints.
@@ 2 String foam and the shower formation
In string models for space-time foam, the in-vacuo propagation of photons is modified as initially inspired from Liouville QG Ellis:1992eh , with their travel times delayed by amounts proportional to $\eta-1$ (c.f. (2)), due to D-brane space-time defects. These defects break Poincaré invariance, and their recoil during interactions with an open string representing a photon breaks local Lorentz invariance. They give space-time a foamy nature so the situations were termed D-foam. The cosmological setup of D-foam is based on braneworlds moving in a higher-dimensional bulk space that contains a “gas” of D0-branes (D-particles) Ellis:2004ay of type I/IIA strings. We consider it to be present in models of phenomenological interest Ellis:1999uh , Ellis:2000sf , Ellis:2004ay , Mavromatos:2005bu , Ellis:2008gg , Mavromatos:2012ii , Li:2021gah , Li:2021tcw , Li:2022sgs , Li:2023wlo , Li:2024crc as even if the theory admits no D-particles one has effective ones Li:2009tt . As our brane roams in the bulk, they cross it randomly, and collide with the photon resided on it and it is not possible in general to capture the process with local effective Lagrangians.
We have previously investigated stochastic effects of the foam in a model of this kind in Li:2022sgs which will violate CPT, and neutrinos may experience a so-induced CPT violation during propagation in this QG “medium”, dictated by the dispersion relations:
@@
In order to discuss cosmic photon-induced showers let us first derive the photon dispersions, via considering the nontrivial momentum transfer $\Delta k_{i}\equiv rk_{i}$, in each collision of a photon off a D-(particle)-defect, which recoils at a velocity, $u$ with its components along the brane, $u_{i}=rg_{s}k_{i}/M_{s}$ Ellis:1999uh , Ellis:2000sf , Ellis:2004ay . We assume isotropic random recoil, as in Li:2022sgs , i.e., the momentum fraction variable $r$ vanishes on average $\langle\!\langle r\rangle\!\rangle=0$ and $\langle\!\langle r^{2}\rangle\!\rangle\eqqcolon\sigma^{2}<1$, where $\langle\!\langle\cdot\cdot\cdot\rangle\!\rangle$’s are statistical averages over a collection of quantum-fluctuating defects met by the photon. Here, the variance $\sigma$ plays a role similar in nature to the parameter $\zeta$ for neutrino species ($\sigma\neq\zeta$ in general). [^1] The topologically nontrivial effects of the foam on particle propagation do not obey the principle of equivalence Ellis:2003ua , Ellis:2003sd , in the sense of being universal for all types of excitations of the Standard Model; moreover for reasons of charge conservation such a medium of defects is “transparent” for charged particles, such as electrons (as indicated by experiment Li:2022ugz ). The metric surrounding the struck defect is of Finsler type $G_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$ and $h_{0i}=u_{i}$ modifying the energy of the photon via $k^{\mu}k^{\nu}G_{\mu\nu}=0$. One obtains after averaging over foam particles the modified energy, $\omega\simeq k[1+\frac{1}{2}\sigma^{2}k^{2}/M_{D}^{2}]$, used in the kinematics: $\omega=\omega^{\prime}+\langle\!\langle M_{D}u^{2}/2\rangle\!\rangle$, where $\frac{M_{s}}{2g_{s}}u^{2}$ is the kinetic energy of heavy defects in the case of nonrelativistic recoil, $u_{i}\ll 1$. The final dispersion relation found from $\omega^{\prime}$ in the leading approximation yields,
@@
here $\sigma^{2}$ a free parameter, depends on foam density $n_{D}$ (and microscopic details that cannot be determined quantitively presently). It is clear that this subluminal dispersion effect leads, through the appropriate group velocity ($v\coloneqq\partial\omega/\partial k$ assumed to hold), to a QG refractive index (2) for photons: $v\simeq 1-\omega\sigma^{2}/M_{D}$, of the kind favored Li:2021gah by current time-lag studies of GRB photons Shao:2009bv , Zhang:2014wpb , Amelino-Camelia:2016ohi , Xu:2016zxi , Xu:2016zsa , Xu:2018ien , Liu:2018qrg , Chen:2019avc , Zhu:2021pml , Zhu:2022usw , Song:2024and , Song:2025myx , Song:2025qej , Song:2025akr . For a uniform D-defect density, at least for late eras of the Universe history where $\sigma$ is constant, an order of magnitude $\sigma^{2}\sim{\cal O}(1)$ is sufficient to reproduce the finding of light-speed variation mentioned earlier, assuming a Planckian D-particle mass $\sim 10^{18}$ GeV. In cases of lower $M_{D}$ the required $\sigma^{2}\ll 1$ is easily satisfied with more dilute foam populations.
@@
It has been argued Satunin:2023yvj however that, due to the presence of the modified photon dispersion (1) which, as we see, may originate from a string-inspired scenario (c.f. (4)) the cross-sections of standard (perturbative) processes are modified. In particular, any allowed interaction channel is suppressed for subluminal photons, since such LV photons lack energy compared to the momentum. Earlier as argued in Vankov:2002gt they may even suppress the cascade generated by cosmic $\gamma$-rays, causing the showers to be formed deeper in the atmosphere and in detectors. The absence of the effect in the measured spectrum sets restrictive (shower formation) constraints on $E_{\textrm{LV}}$ Satunin:2023yvj . However, these considerations also concern strict energy–momentum conservation (assumed in prior works), not only modified dispersion relation.
@@
In general, in the space-time foam picture energy is conserved only in a statistical sense (see the review Sarkar:2002mg ), however, even this can be violated due to unobserved degrees of freedom (from the point of view of a low-energy braneworld observer) associated with the D-particle recoil in the model of Li:2022sgs . Indeed, there are quantum fluctuations $\delta E_{D}$ in the total energy in particle reactions Ellis:2000sf . For a three-body decay process in the recoil foam adapted to the situation here discussed, the energy of the incoming particle $E$ fluctuates, due to the presence of stochastic D-foam, by the amount of $(\varsigma_{I}/M_{D})p^{2}$ where $p$, the momentum of the particle emitted during the decay of a particle of effective energy $E+\delta E_{D}$. The proportional factor $\varsigma_{I}$ expressing the intensity of such fluctuations can vary independently of the foam dispersion factor $\sigma$ and values $\varsigma_{I}>0$, as in principle permitted by the theory, will be seriously considered here. Now we are going to see how this affects the conclusion drawn from other LV approaches Vankov:2002gt , Myers:2003fd , yielding unacceptable anomalies in the formation of cosmic $\gamma$-ray showers. As we shall see, the peculiarity of this foam in resulting in such energy violations is crucial for accommodating the result of light-speed variation, without conflicting with the constraints from shower measurements.
@@
At the Earth photons initiate almost purely EM showers via pair creation and bremsstrahlung Risse:2007sd . Additional processes are important only at highest energy ($10^{19}-10^{20}$ eV) so for GeV to sub-PeV $\gamma$-rays we ignore them for simplicity except where noted. Lacking at present a complete theory of the matter/D-foam interactions we do not deal with the cross-sections directly and shall restrict ourselves to calculate the length $\lambda_{f}$, over which the processes are considered taking place according to uncertainty principle, that is formation (or coherent) length Klein:1998du .
@@
where $p$, $p^{\prime}$ indicate the outgoing momenta of electron and positron and photon’s motion is in $z$ axis, with its momentum $k$. The nucleus’ Coulomb field has a larger component perpendicular to $z$ axis, leading to a transverse momentum kick $\Delta q_{\perp}\sim{\cal O}(m_{e})$ ($m_{e}$, the electron rest mass), while the longitudinal momentum $\Delta q_{z}$ is small, $\Delta q_{z}\ll k,\Delta q_{\perp}$ Rubtsov:2012kb , due to the photon’s high energy $\omega\gg 2m_{e}$. For this regime one can ignore collision angles. In the presence of string D-foam, one should make use of the dispersion including the foam induced LV (4). For electron (positron) its law of dispersion is preserved as noted, for gauge invariance Ellis:2003ua , Ellis:2003sd , so we have $E^{2}-m_{e}^{2}=p^{2}$ and then,
@@
We come to discuss the formation length, $\lambda_{f}$, dictated by Heisenberg’s uncertainty principle. While the latter suffers a modification, due to the existence of the minimal (string) length $\ell_{s}=1/M_{s}$ in target space-time Veneziano:1986zf ,
@@
where $\langle X_{0}\rangle\simeq 50~\textrm{g}\cdot\textrm{cm}^{-2}$ with $\sigma_{\textrm{BH}}\approx 0.51$ b in air. Then, the probability for a $\gamma$-ray to develop a shower and then be detected as penetrating through the Earth’s atmosphere is given in Satunin:2023yvj and when adapted to our case reads
@@
It is from these considerations that interesting constraints on subluminal-photon linear LV were derived recently from the LHAASO data for TeV–PeV photon-induced air showers LHAASO:2021gok . Unlike the linear LV model of Vankov:2002gt , Myers:2003fd , however, in our case, such limits are not imposed on dispersion parameter $\sigma^{2}$ which is related to $E_{\textrm{LV}}$ in Eq. (1) but for $2\sigma^{2}-\varsigma_{I}$. Naively, one could use the result they derive to make an order of magnitude estimate in the parameters of the present model: [^2] We notice that for the result of Satunin:2023yvj , the effect of LV-modified pair creation on soft cosmic background photons leading to the increase of the flux is also taken into account, whilst it is subleading compared to the shower formation effects (also for priorly studied spectra Satunin:2019gsl , Rubtsov:2016bea ). However, the existence of such modifications in the present model has not been clarified, hence temporally ignored here and is left for future investigation.
@@
The other effect that opposes those from general and/or field-theoretic description of subluminal LV photons is that if $\varsigma_{I}>2\sigma^{2}$, Eq. (2) turns into enhancement which implies that, the formation length increases instead of being shortened. And the factor $S$ even becomes infinite at the critical energy $\omega_{\textrm{cr}}=\bigl[\frac{8M_{D}m_{e}^{2}}{(\varsigma_{I}-2\sigma^{2})}\bigr]^{1/3}$ and negative above it, while the situation could be compensated by suppression factors associated with interactions that can disrupt coherence over such a longer length. The effects, that increase $\Delta q_{z}$ and reduce $\lambda_{f}$, include, for example, the Landau–Pomeranchuk–Migdal effect (see, e.g., Klein:1998du , Risse:2007sd ) due to multiple scattering centers. The so-reduced BH cross-section delays the development of the EM cascade initiated by a photon competing with foam effects which decrease now $\langle X_{0}\rangle$ and $\langle X_{\max}\rangle$. It is worth noting that the length of shower development $\Delta X$ shifting $X_{\max}$ as $X_{\max}=X_{0}+\Delta X$ is assumed unmodified in the above discussion, since the secondary interactions in the cascade are less energetic than the primary one. While, in principle, it may be affected by “exotic” physics studied here, not accounting for this leaves the estimate (19) more conservative.
@@
where $p$ and $p^{\prime}$ stand now for the electron momenta before and after the interaction, and the photon emission angle is neglected for highly relativistic electron $E\gg m_{e}$. One difference herein is that, unlike pair creation where the energy partition in the produced $ee^{+}$ pair is most likely symmetric as we assumed above, [^3] It was shown Rubtsov:2012kb that LV may enhance the leading particle effect as the cross-section is peaked at $E\sim\omega$ and $\omega-E\sim\omega$. But, there is no good reason to expect the same for D-foam cases where, unlike the model studied therein, to first order in $1/M_{s}$, the matrix element for any process agrees with the one in the special-relativistic field theory so we consider symmetric pairs as an illustration. the probability of asymmetric share is significant in bremsstrahlung. The resulting spectrum is dominated by photons emitted at very low energy ($\omega\ll E$) and diverges as $\omega\rightarrow 0$ ($1/\omega$ dependence). Then the effects of foam for such configurations of soft bremsstrahlung photons are negligible, compared to the first two conventional, leading terms as in (8):
@@
While for photons at moderate energies, the process may be mildly affected, nevertheless it is important to note that in order for an electron to emit a real photon as it interacts with the field of the nucleus, part of its surrounding virtual photon field shakes loose, a case for which no modification is induced in this scenario. As we have explained, LV effect within such D-brane models is due to multiple interactions between photons and D-particles, hence cannot be relevant in any mere process of photon production. This is another difference from the phenomenological consideration of Vankov:2002gt , where due to electron LV effect, bremsstrahlung is strongly suppressed, similarly to pair creation on nuclei by photons, reducing the formation length $\lambda_{f}$ from (21), and hence the probability for radiation. There is currently no direct limit from bremsstrahlung though as the latter is usually viewed as unmodified, for practical (LV constraining) purposes, as also mentioned before.